OpenStudy (anonymous):

Which table shows a proportional relationship between x and y? (posted below)

OpenStudy (anonymous):

I think it is c, is that right?

OpenStudy (welshfella):

check out the y values and see if y = kx for each entry k will be some constant like 2,3 etc

OpenStudy (welshfella):

No because of the 0

OpenStudy (anonymous):

y because of the zero?

OpenStudy (welshfella):

well you have 8 = 4*2 as the second entry - making k = 4 but for first entry you have 8 = k * 0 which of course is wrong

OpenStudy (welshfella):

the values of k must be the same for all entries in the table

OpenStudy (welshfella):

divide the values of y by the values of x and see what you get try table B

OpenStudy (anonymous):

ok so b was my second answer is it proportional ?

OpenStudy (welshfella):

do you get it Table B 3 / 1 = ? 6 / 2 = ? then do third and 4th entry

OpenStudy (welshfella):

see for yourself - do the divisions

OpenStudy (welshfella):

3/1 = 3 6/2 = 3 12/4 = ? 15/5 = ?

OpenStudy (welshfella):

6 /2 = 3 not 12

OpenStudy (anonymous):

I know I was just about to correct it

OpenStudy (welshfella):

ok

OpenStudy (anonymous):

12/4= 3 and... 15/5= 3

OpenStudy (welshfella):

right so ity must be B

OpenStudy (anonymous):

oh! I didn't know how to solve these equations and now I do! Thank You :D

OpenStudy (anonymous):

oh and I believe number two is proportional and the ratio of dollars is 50$ per month is that right?

OpenStudy (anonymous):

just to check